The hypotenuse of a right triangle is 13 cm. One leg of the triangle is 7cm longer than the other. What are the lengths of the two legs?
Answer by stanbon(48535)

The hypotenuse of a right triangle is 13 cm. One leg of the triangle is 7cm longer than the other. What are the lengths of the two legs?
-----------------------------------------
Let one leg be "x"
The other leg will be "x+7"
------------------
Equation:
x^2 + (x+7)^2 = 13^2
2x^2 + 14x + 49 - 169 = 0
2x^2 + 14x -120 = 0
x^2 + 7x - 60 = 0
x^2 + 12x - 5x - 60 = 0
x(x+12) -5(x+12) = 0
(x+12)(x-5) = 0
Positive solution:
x = 5 (one leg)
x+7 = 12 (other leg)
========================
Cheers,
Stan H.